The following table shows the total output per hour produced in a factory at various levels of employment of labor. The firm sells each unit of output at $2 and each worker is paid a wage of $12. What is the optimal number of workers (where wage equals marginal revenue)?

Number of Workers Total Output Marginal Output
1 8
2 18 10
3 30 12
4 41 11
5 50 9
6 56 6
7 60 4
8 61 1
9 62 -1

Answers:

A. 4

B. 8

C. 5

D. 7

E. 6
Answer #1

E. 6

At optimum marginal output = W/P

W/P = 12/2= 6

Marginal output is 6 when number of workers is 6